Evaluate the integral of 1/y² Cos (1/y) dy 


1
Expert's answer
2021-03-10T07:08:18-0500

Evaluate the integral

"\displaystyle\int\dfrac{1}{y^2}\cos \left(\frac{1}{y}\right) dy"


Note that


"d\left(\dfrac{1}{y}\right)=\left(\dfrac{1}{y}\right)'dy=-\dfrac{1}{y^2}dy"


Then


"\displaystyle\int\dfrac{1}{y^2}\cdot\cos \left(\frac{1}{y}\right) dy ="


"\displaystyle\int\cos \left(\frac{1}{y}\right)\cdot\dfrac{1}{y^2} dy ="


"-\displaystyle\int\cos \left(\frac{1}{y}\right)\cdot\left(-\dfrac{1}{y^2}\right) dy ="


"-\displaystyle\int\cos \left(\frac{1}{y}\right)d\left(\dfrac{1}{y}\right) ="


"-\sin \left(\dfrac{1}{y}\right)+C"